# Large Language Model-Brained GUI Agents: A Survey

### v6

@@ 1.3 Survey Structure
Innovations: A Roadmap, tier2] [5.7 Advanced Enhancements, tier2 [5.7.6 Summary & Takeaways, tier3] [5.7.5 Reinforcement Learning, tier3] [5.7.4 Self-Evolution, tier3] [5.7.3 Self-Reflection, tier3] [5.7.2 Multi-Agent Framework, tier3] [5.7.1 Computer Vision-Based GUI Parsing, tier3, yshift=-90pt] ] [5.6 Memory, tier2 [5.6.2 Long-Term Memory, tier3] [5.6.1 Short-Term Memory, tier3] ] [5.5 Actions Execution, tier2 [5.5.4 Summary, tier3] [5.5.3 AI Tools, tier3] [5.5.2 Native API Calls, tier3] [5.5.1 UI Operations, tier3] ] [5.4 Model Inference, tier2 [5.4.3 Complementary Outputs, tier3] [5.4.2 Action Inference, tier3] [5.4.1 Planning, tier3] ] [5.3 Prompt Engineering, tier2 ] [5.2 Operating Environment, tier2 [5.2.3 Environment Feedback, tier3] [5.2.2 Environment State Perception, tier3] [5.2.1 Platform, tier3, yshift=-40pt] ] [5.1 Architecture and Workflow in a Nutshell, tier2] ] [4 Evolution and Progression of
@@ 5.7 Advanced Enhancements
5.7.1 Computer Vision-Based GUI Parsing
@@
CV models offer a non-intrusive solution for parsing GUIs directly from screenshots, enabling the detection, localization, segmentation, and even functional estimation of widgets [194, 91, 195, 196]. This approach allows agents to interpret the visual structure and elements of the GUI without relying on system-level tools or internal metadata, which may be unavailable or incomplete. CV-based GUI parsing provides agents with valuable insights into interactive components, screen layout, and widget functionalities based solely on visual cues, enhancing their ability to recognize and act upon elements on the screen. Figure 9 provides an illustrative example of how a CV-based GUI parser works. While standard API-based detection captures predefined widgets, the CV model can identify additional elements, such as thumbnails and canvases, which may not have explicit API representations in the PowerPoint interface. This enhances widget recognition, allowing the agent to detect components beyond the scope of API detection.
@@
Such CV-based GUI parsing layers provide critical grounding information that significantly enhances an agent’s ability to interact accurately and intuitively with diverse GUIs. This is particularly beneficial for handling custom or non-standard elements that deviate from typical accessibility protocols. Additionally, prompting methods like iterative narrowing have shown promise in improving the widget grounding capabilities of VLMs [197]. Together, these approaches pave the way for more adaptable and resilient GUI agents, capable of operating effectively across a broader range of screen environments and application contexts.
@@ 5.8 From Foundations to Innovations: A Roadmap
FoundationstoInnovations Frameworks Web [16, 186] [216, 217, 218, 219, 220, 221, 222, 223, 224, 225, 226, 227, 228, 229, 230, 231, 232, 233, 234, 235, 236, 237, 238, 239] Mobile [17, 200, 146] [144, 240, 238] [232, 233, 239] [241, 242, 243, 244, 245, 246, 247, 248, 249, 250, 251, 252] Computer [18, 149, 150] [200, 246, 233] [253, 254, 255, 251] [238, 239] Cross- Platform [232, 238] [233, 200] [251, 256] [239] Data Web [257, 258, 259, 260, 261, 262, 263, 264, 265, 266] Mobile [170, 132] [240, 260] [263, 261] [267, 268, 269, 270, 271, 272, 273, 274, 275, 276, 277, 278, 279, 280, 281, 266] Computer [170, 266] [282, 280, 281, 283] Cross- Platform [280, 266] [170, 263] [261, 264] [260, 281] Models Foundation Models [80, 81, 151] [284, 285, 286, 287, 288, 289, 290, 291, 292, 293, 294] Web [24, 15, 170] [141, 231, 264] [266, 295, 296, 297, 298, 299, 300, 301] Mobile [24, 15, 170] [214, 267, 277] [275, 302, 274, 303, 304, 305, 306, 147, 307, 308, 281, 299, 300, 301] [266] Computer [24, 15, 282] [309, 308, 281] [170, 300, 266] [301, 264, 283] Cross- Platform [15, 24, 266] [170, 281, 264] [308, 299, 300, 301] Evaluation Web [228, 220, 24] [130, 133, 134] [257, 258, 259, 260] [310, 311, 312, 313, 262, 314, 315, 316, 317, 318, 319, 320, 321, 322, 323, 324, 325, 326, 327] Mobile [278, 326] [328, 144, 132] [213, 146, 329, 330, 331, 271, 332, 333, 334, 335, 260, 24, 336, 337] Computer [24, 328] [326, 97, 338] [317, 318, 319] [339, 340, 341] Cross- Platform [254] [260, 318, 319] [328, 24, 326] Applications GUI Testing [342, 113] [343, 344, 345, 346, 347, 348, 349, 350, 351, 352] Virtual Assistants [120] [353, 354, 355, 356, 357, 358, 359, 360, 361, 362, 363, 364, 365]
@@
TABLE VI: Overview of LLM-brained GUI agent frameworks (Part I).
@@
TABLE VII: Overview of LLM-brained GUI agent frameworks (Part II).
Agent Platform Perception Action Model Architecture Highlight Link
@@
TABLE VIII: Overview of LLM-brained GUI agent frameworks (Part III).
Agent Platform Perception Action Model Architecture Highlight Link
TABLE IX: Overview of LLM-brained GUI agent frameworks (Part IV).
Agent Platform Perception Action Model Architecture Highlight Link
6 LLM-Brained GUI Agent Framework
To provide a comprehensive view of the field, we first summarize key frameworks across web, mobile, computer and cross-platform in Tables VI, VII, VIII, and IX, highlighting their features, target platforms, and notable innovations. These summaries serve as an entry point to understanding the breadth of development in LLM-brained GUI agents. Following this overview, we offer a detailed discussion of each framework, examining their foundational design principles, technical advancements, and the specific challenges they address in the realm of GUI automation. By delving into these aspects, we aim to provide deeper insights into how these agents are shaping the future of human-computer interaction and task automation, and the critical role they play in advancing this transformative field.
@@ 6.1 Web GUI Agents
Advancements in web GUI agents have led to significant strides in automating complex tasks within diverse and dynamic web environments. Recent frameworks have introduced innovative approaches that leverage multimodal inputs, predictive modeling, and task-specific optimizations to enhance performance, adaptability, and efficiency. In this subsection, we delve into these frameworks, highlighting their unique contributions and how they collectively push the boundaries of web-based GUI automation.
@@ 6.2 Mobile GUI Agents
The evolution of mobile GUI agents has been marked by significant advancements, leveraging multimodal models, complex architectures, and adaptive planning to address the unique challenges of mobile environments. These agents have progressed from basic interaction capabilities to sophisticated systems capable of dynamic, context-aware operations across diverse mobile applications.
@@
Addressing the need for efficient handling of multi-step tasks with lower computational costs, AutoDroid [144] combines LLM-based comprehension with app-specific knowledge. Using an HTML-style GUI representation and a memory-based approach, AutoDroid reduces dependency on extensive LLM queries. Its hybrid architecture of cloud and on-device models enhances responsiveness and accessibility, making AutoDroid a practical solution for diverse mobile tasks.
@@ 6.3 Computer GUI Agents
Computer GUI agents have evolved to offer complex automation capabilities across diverse operating systems, addressing challenges such as cross-application interaction, task generalization, and high-level task planning.
@@
Advancements in computer GUI agents have led to the development of sophisticated frameworks capable of handling complex tasks across desktop environments. These agents have evolved from simple automation tools to intelligent systems that leverage multimodal inputs, advanced architectures, and adaptive learning to perform multi-application tasks with high efficiency and adaptability.
@@ 6.4 Cross-Platform GUI Agents
Cross-platform GUI agents have emerged as versatile solutions capable of interacting with various environments, from desktop and mobile platforms to more complex systems. These frameworks prioritize adaptability and efficiency, leveraging both lightweight models and multi-agent architectures to enhance cross-platform operability. In this subsection, we explore key frameworks that exemplify the advancements in cross-platform GUI automation.
@@ 7 Data for Optimizing LLM-Brained GUI Agents
As GUI agents continue to gain traction, researchers have focused on assembling datasets that represent a broad spectrum of platforms and capture the diverse intricacies of GUI environments. These datasets are pivotal in training models that can generalize effectively, thanks to their coverage of varied interfaces, workflows, and user interactions. To ensure comprehensive representation, innovative methodologies have been employed to collect and structure these data assets. In the sections that follow, we detail an end-to-end pipeline for data collection and processing tailored to training GUI-specific LLMs. We also examine significant datasets from various platforms, providing insights into their unique features, the methodologies used in their creation, and their potential applications in advancing the field of LLM-brained GUI agents. To provide a structured overview, Tables X,  XI, and XII summarize these datasets across differernt platforms, highlighting their key attributes and contributions to the evolution of GUI agent research.

### v7

@@ 1.3 Survey Structure
Innovations: A Roadmap, tier2] [5.7 Advanced Enhancements, tier2 [5.7.6 Summary & Takeaways, tier3] [5.7.5 Reinforcement Learning, tier3] [5.7.4 Self-Evolution, tier3] [5.7.3 Self-Reflection, tier3] [5.7.2 Multi-Agent Framework, tier3] [5.7.1 Computer Vision-Based GUI Grounding, tier3, yshift=-90pt] ] [5.6 Memory, tier2 [5.6.2 Long-Term Memory, tier3] [5.6.1 Short-Term Memory, tier3] ] [5.5 Actions Execution, tier2 [5.5.4 Summary, tier3] [5.5.3 AI Tools, tier3] [5.5.2 Native API Calls, tier3] [5.5.1 UI Operations, tier3] ] [5.4 Model Inference, tier2 [5.4.3 Complementary Outputs, tier3] [5.4.2 Action Inference, tier3] [5.4.1 Planning, tier3] ] [5.3 Prompt Engineering, tier2 ] [5.2 Operating Environment, tier2 [5.2.3 Environment Feedback, tier3] [5.2.2 Environment State Perception, tier3] [5.2.1 Platform, tier3, yshift=-40pt] ] [5.1 Architecture and Workflow in a Nutshell, tier2] ] [4 Evolution and Progression of
@@ 2 Related Work
Li et al., [54] A survey on benchmarking evaluations, applications, and challenges of visual LLMs.
@@
Dang et al., [61] A Survey on GUI agents.
Liu et al., [62] A Survey on GUI agent on phone automation.
Hu et al., [63] A Survey on MLLM based agents for OS.
@@ 5.7 Advanced Enhancements
TABLE VI: A summary of of GUI grounding models and benchmarks.
Model/ Benchmark Platform Foundation Model Size Dataset Input Output Highlight Link
OmniParser [174] Mobile, Desktop, and Web BLIP-2 [197], YOLOv8 [198] / 67,000 UI screenshots with bounding box annotations and 7,185 icon–description pairs generated using GPT-4 UI screenshots IDs, bounding boxes, and descriptions of interactable elements Introduces a purely vision-based screen parsing framework for general UI understanding without external information, significantly improving action prediction accuracy for LLM-driven agents https://github.com/microsoft/OmniParser
Iterative Narrowing [199] Mobile, Web, and Desktop Qwen2-VL and OS-Atlas-Base / ScreenSpot [24] A GUI screenshot and a natural language query (x,y) coordinates representing the target location in the GUI Progressively crops regions of the GUI to refine predictions, enhancing precision for GUI grounding tasks https://github.com/ant-8/GUI-Grounding-via-Iterative-Narrowing
Iris [200] Mobile (iOS, Android), Desktop (Windows, macOS), and Web Qwen-VL [201] 9.6B 850K GUI-specific annotations and 150K vision–language instructions High-resolution GUI screenshots with natural language instructions Referring: Generates detailed descriptions of UI elements. Grounding: Locates UI elements on the screen. Information-Sensitive Cropping for efficient handling of high-resolution GUI images, and Self-Refining Dual Learning to iteratively enhance GUI grounding and referring tasks without additional annotations /
Attention-driven Grounding [202] Mobile, Web, and Desktop MiniCPM-Llama3-V 2.5 8.5B Mind2Web [203], ScreenSpot [24], VisualWebBench [204] GUI screenshots and textual user queries Element localization via bounding boxes, text-to-image mapping for grounding, and actionable descriptions of GUI components Utilizes attention mechanisms in pre-trained MLLMs without fine-tuning https://github.com/HeimingX/TAG
Aria-UI [205] Web, Desktop, and Mobile Aria [206] 3.9B 3.9 million elements and 11.5 million samples GUI screenshots, user instructions, and action histories Pixel coordinates for GUI elements and corresponding actions A purely vision-based approach avoiding reliance on AXTree-like inputs https://ariaui.github.io
ScreenSpot [24] iOS, Android, macOS, and Windows (benchmark) (benchmark) Over 600 screenshots and 1,200 instructions GUI screenshots accompanied by user instructions Bounding boxes or coordinates of actionable GUI elements A realistic and diverse GUI grounding benchmark covering multiple platforms and a variety of elements https://github.com/njucckevin/SeeClick
ScreenSpot-Pro [207] Windows, macOS, and Linux (benchmark) (benchmark) 1,581 instruction–screenshot pairs covering 23 applications across 5 industries and 3 operating systems High-resolution GUI screenshots paired with natural language instructions Bounding boxes for locating target UI elements Introduces a high-resolution benchmark for professional environments https://github.com/likaixin2000/ScreenSpot-Pro-GUI-Grounding
5.7.1 Computer Vision-Based GUI Grounding
@@
CV models offer a non-intrusive solution for GUI grounding directly from screenshots, enabling the detection, localization, segmentation, and even functional estimation of widgets [209, 95, 210, 211]. This approach allows agents to interpret the visual structure and elements of the GUI without relying on system-level tools or internal metadata, which may be unavailable or incomplete. CV-based GUI parsing provides agents with valuable insights into interactive components, screen layout, and widget functionalities based solely on visual cues, enhancing their ability to recognize and act upon elements on the screen. Figure 9 provides an illustrative example of how a CV-based GUI parser works. While standard API-based detection captures predefined widgets, the CV model can identify additional elements, such as thumbnails and canvases, which may not have explicit API representations in the PowerPoint interface. This enhances widget recognition, allowing the agent to detect components beyond the scope of API detection. We show an overview of related GUI grounding models and benchmarks in Table VI.
@@
Such CV-based GUI grounding layers provide critical grounding information that significantly enhances an agent’s ability to interact accurately and intuitively with diverse GUIs. This is particularly beneficial for handling custom or non-standard elements that deviate from typical accessibility protocols. Additionally, prompting methods like iterative narrowing have shown promise in improving the widget grounding capabilities of VLMs [199]. Together, these approaches pave the way for more adaptable and resilient GUI agents, capable of operating effectively across a broader range of screen environments and application contexts.
Several works have introduced benchmarks to evaluate the GUI grounding capabilities of models and agents. For instance, ScreenSpot [24] serves as a pioneering benchmark designed to assess the GUI grounding performance of LLM-powered agents across diverse platforms, including iOS, Android, macOS, Windows, and web environments. It features a dataset with over 600 screenshots and 1,200 instructions, focusing on complex GUI components such as widgets and icons. This benchmark emphasizes the importance of GUI grounding in enhancing downstream tasks like web automation and mobile UI interaction. Building upon this, ScreenSpot-Pro [207] extends the scope to more professional, high-resolution environments. This evolved version includes 1,581 tasks with high-quality annotations, encompassing domains such as software development, creative tools, CAD, scientific applications, and office productivity. Key features of ScreenSpot-Pro include authentic high-resolution screenshots and meticulous annotations provided by domain experts.
These benchmarks provide critical evaluation criteria for assessing GUI grounding capabilities, thereby advancing the development of GUI agents for improved GUI understanding and interaction.
@@ 5.8 From Foundations to Innovations: A Roadmap
FoundationstoInnovations Frameworks Web [16, 190] [230, 231, 232, 233, 234, 235, 236, 237, 238, 239, 240, 241, 242, 243, 244, 245, 246, 247, 248, 249, 250, 251, 252, 253, 254, 255, 256] Mobile [17, 214, 150] [148, 257, 252] [246, 247, 253] [256, 258, 259, 260, 261, 262, 263, 264, 265, 266, 267, 268, 269, 270] Computer [256] [18, 153, 154] [214, 263, 247] [271, 272, 273, 269] [252, 253, 274] Cross- Platform [246, 252] [247, 214] [269, 275] [253, 256] Data Web [203] [276, 277, 278, 279, 280, 281, 282, 283, 284, 285] Mobile [174, 136] [257, 278] [281, 279] [286, 287, 288, 289, 290, 291, 292, 293, 294, 295, 296, 297, 298, 299, 300, 284, 285] Computer [174, 284] [301, 299, 300, 302] Cross- Platform [285] [299, 284] [174, 281] [279, 282] [278, 300] Models Foundation Models [197, 201] [84, 85, 155] [303, 304, 305, 306, 307, 308, 309, 310, 311] Web [24, 15, 174] [145, 245, 282] [284, 312, 313, 314, 315, 316, 317, 318] Mobile [24, 15, 174] [228, 286, 296] [294, 319, 293, 320, 321, 322, 323, 151, 324, 325, 300, 316, 317, 318] [284] Computer [24, 15, 301] [326, 325, 300] [174, 317, 284] [318, 282, 302] Cross- Platform [15, 24, 284] [174, 300, 282] [325, 316, 317, 318] Evaluation Web [204] [242, 234, 24] [134, 137, 138] [203, 276, 277, 278] [327, 328, 329, 330, 280, 331, 332, 333, 334, 335, 336, 337, 338, 339, 340, 341, 342, 343, 344, 345] Mobile [297, 342] [346, 148, 136] [227, 150, 347, 348, 349, 290, 350, 351, 352, 353, 278, 24, 354, 355, 356, 357, 358] Computer [24, 346] [342, 101, 359] [334, 335, 336] [360, 361, 362] Cross- Platform [272] [278, 335, 336] [346, 24, 342] Applications GUI Testing [363, 117] [364, 365, 366, 367, 368, 369, 370, 371, 372, 373] Virtual Assistants [124] [374, 375, 376, 377, 378, 379, 380, 381, 382, 383, 384, 385, 386]
@@
6 LLM-Brained GUI Agent Framework
We offer a detailed discussion of each framework, examining their foundational design principles, technical advancements, and the specific challenges they address in the realm of GUI automation. By delving into these aspects, we aim to provide deeper insights into how these agents are shaping the future of human-computer interaction and task automation, and the critical role they play in advancing this transformative field.
TABLE VII: Overview of LLM-brained GUI agent frameworks on web platforms (Part I).
@@ 6 LLM-Brained GUI Agent Framework
TABLE VIII: Overview of LLM-brained GUI agent frameworks on web platforms (Part II).
Agent Platform Perception Action Model Architecture Highlight Link
@@
PAE [254] Web Screenshots Standard UI Operations Claude 3 Sonnet [155], Qwen2VL-7B [304], and LLaVa-1.6 [309] A multi-agent architecture involving a task proposer to suggest tasks, an agent policy to perform tasks, and an autonomous evaluator to assess success and provide feedback using RL. Autonomous skill discovery in real-world environments using task proposers and reward-based evaluation https://yanqval.github.io/PAE/
@@
WEPO [255] Web HTML and DOM Standard UI Operations Llama3-8B [83], Mistral-7B [390], and Gemma-2B [391] Single-agent architecture. Incorporates a distance-based sampling mechanism tailored to the DOM tree structure, enhancing preference learning by distinguishing between salient and non-salient web elements with DPO [392]. /
@@ 6.1 Web GUI Agents
Advancements in web GUI agents have led to significant strides in automating complex tasks within diverse and dynamic web environments. Recent frameworks have introduced innovative approaches that leverage multimodal inputs, predictive modeling, and task-specific optimizations to enhance performance, adaptability, and efficiency. In this subsection, we first summarize key web GUI agent frameworks in Tables VII and VIII, then delve into representative frameworks, highlighting their unique contributions and how they collectively push the boundaries of web-based GUI automation.
@@
TABLE IX: Overview of LLM-brained GUI agent frameworks on mobile platforms (Part I).
Agent Platform Perception Action Model Architecture Highlight Link
TABLE X: Overview of LLM-brained GUI agent frameworks on mobile platforms (Part II).
Agent Platform Perception Action Model Architecture Highlight Link
AutoDroid-V2 [264] Android Mobile Devices Structured GUI Representations Multi-step scripts of standard UI operations and API calls Llama-3.1-8B [83] Script-based architecture. Converts GUI task automation into a script generation problem, enhancing efficiency and task success rates. /
@@ 6.2 Mobile GUI Agents
The evolution of mobile GUI agents has been marked by significant advancements, leveraging multimodal models, complex architectures, and adaptive planning to address the unique challenges of mobile environments. These agents have progressed from basic interaction capabilities to sophisticated systems capable of dynamic, context-aware operations across diverse mobile applications. We first provide an overview of mobile GUI agent frameworks in Tables IX and X.
@@
Addressing the need for efficient handling of multi-step tasks with lower computational costs, AutoDroid [148] combines LLM-based comprehension with app-specific knowledge. Using an HTML-style GUI representation and a memory-based approach, AutoDroid reduces dependency on extensive LLM queries. Its hybrid architecture of cloud and on-device models enhances responsiveness and accessibility, making AutoDroid a practical solution for diverse mobile tasks. AutoDroid-V2 [264] enhances its predecessor AutoDroid, by utilizing on-device language models to generate and execute multi-step scripts for user task automation. By transforming dynamic and complex GUI elements of mobile apps into structured app documents, it achieves efficient and accurate automation without depending on cloud-based resources. The script-based approach reduces computational overhead by minimizing query frequency, thereby improving task efficiency and addressing the limitations of stepwise agents. This advancement enables privacy-preserving and scalable task automation on mobile platforms.
@@
TABLE XI: Overview of LLM-brained GUI agent frameworks on computer platforms.
Agent Platform Perception Action Model Architecture Highlight Link
PC Agent [274] Windows Computer Screenshots and event-based tracking Standard UI Operations Qwen2-VL-72B-Instruct [304] and Molmo [398] A planning agent for decision-making combined with a grounding agent for executing actions. Human cognition transfer framework, which transforms raw interaction data into cognitive trajectories to enable complex computer tasks. https://gair-nlp.github.io/PC-Agent/
@@ 6.3 Computer GUI Agents
Computer GUI agents have evolved to offer complex automation capabilities across diverse operating systems, addressing challenges such as cross-application interaction, task generalization, and high-level task planning. They have led to the development of sophisticated frameworks capable of handling complex tasks across desktop environments. These agents have evolved from simple automation tools to intelligent systems that leverage multimodal inputs, advanced architectures, and adaptive learning to perform multi-application tasks with high efficiency and adaptability. We provide an overview of computer GUI agent frameworks in Table XI.
@@
TABLE XII: Overview of LLM-brained cross-platform GUI agent frameworks.
Agent Platform Perception Action Model Architecture Highlight Link
InfiGUIAgent [256] Mobile, Web, Desktop Raw screenshots Standard UI operations. Qwen2-VL-2B [304] Single-agent architecture enhanced by hierarchical reasoning. Introduces native reasoning skills, such as hierarchical and expectation-reflection reasoning, enabling advanced and adaptive task handling. https://github.com/Reallm-Labs/InfiGUIAgent
@@ 6.4 Cross-Platform GUI Agents
Cross-platform GUI agents have emerged as versatile solutions capable of interacting with various environments, from desktop and mobile platforms to more complex systems. These frameworks prioritize adaptability and efficiency, leveraging both lightweight models and multi-agent architectures to enhance cross-platform operability. In this subsection, we first We overview cross-platform GUI agent frameworks in Table XII, then explore key frameworks that exemplify the advancements in cross-platform GUI automation.
@@ 7 Data for Optimizing LLM-Brained GUI Agents
As GUI agents continue to gain traction, researchers have focused on assembling datasets that represent a broad spectrum of platforms and capture the diverse intricacies of GUI environments. These datasets are pivotal in training models that can generalize effectively, thanks to their coverage of varied interfaces, workflows, and user interactions. To ensure comprehensive representation, innovative methodologies have been employed to collect and structure these data assets. In the sections that follow, we detail an end-to-end pipeline for data collection and processing tailored to training GUI-specific LLMs. We also examine significant datasets from various platforms, providing insights into their unique features, the methodologies used in their creation, and their potential applications in advancing the field of LLM-brained GUI agents.